# Blockchain Address Poisoning

### v1

@@ Abstract
In many blockchains, e.g., Ethereum, Binance Smart Chain (BSC), the primary representation used for wallet addresses is a hardly memorable 40-digit hexadecimal string. As a result, users often select addresses from their recent transaction history, which enables blockchain address poisoning. The adversary first generates lookalike addresses similar to one with which the victim has previously interacted, and then engages with the victim to “poison” their transaction history. The goal is to have the victim mistakenly send tokens to the lookalike address, as opposed to the intended recipient. Compared to contemporary studies, this paper provides four notable contributions. First, we develop a detection system and perform measurements over two years on Ethereum and BSC. We identify 13 times the number of attack attempts reported previously—totaling 270M on-chain attacks targeting 17M victims. 6,633 incidents have caused at least 83.8M USD in losses, which makes blockchain address poisoning one of the largest cryptocurrency phishing schemes observed in the wild. Second, we analyze a few large attack entities using improved clustering techniques, and model attacker profitability and competition. Third, we reveal attack strategies—targeted populations, success conditions (address similarity, timing), and cross-chain attacks. Fourth, we mathematically define and simulate the lookalike address-generation process across various software- and hardware-based implementations, and identify a large-scale attacker group that appears to use GPUs. We also discuss defensive countermeasures.
@@ 1 Introduction
Figure 1: Poisoning transfers
@@
We design a detection system and run it between July 1, 2022, and June 30, 2024, on Ethereum and Binance Smart Chain (BSC). Compared to contemporary studies [46, 10], our detection 1) covers all types of poisoning attacks, 2) demonstrates greater robustness against attacker manipulation, and 3) generalizes better across time intervals and chains.
Combining the results from both chains, we identify over 270M attack attempts (i.e., 13 times the previous efforts: 21M [46] and 14M [10]) that target over 17M victims. Out of those, attackers successfully received 6,633 transfers, adding up to over 83.8M USD in ill-gotten gains. This attack appears to be one of the largest cryptocurrency phishing schemes observed in the wild. We validate our results by 1) manually verifying the phished cases, 2) directly comparing the results with the related works, 3) checking the smart contracts used for those attacks, and 4) comparing the data collection results with different parameters (e.g., a window size).
Of independent interest, we leverage our detection system to capture accidental transfers (e.g., typing mistakes in the recipient’s address). We confirm 681 accidental transfers with 5.5M USD losses, based solely on human mistakes which underscores the need for more forgiving user interfaces.
Attackers typically optimize their operations by bundling several transfers into a single transaction and reusing both addresses and contracts. This behavior allows us to link attack addresses through the “guilt-by-association” heuristic [43]. After removing transactions from unrelated bots that occasionally copy poisoning transactions from attackers, we can establish the existence of a few large distinct attack groups which we use to model the economics of address poisoning. For each group, we calculate revenues (i.e., a sum of phished amounts) and costs (i.e., total attack transaction costs). Despite high variability, most large groups manage to stay profitable by succeeding often, while smaller groups sometimes suffer from loss (by losing to the competition). There is no apparent early mover advantage.
@@
Finally, we show that attackers are not as efficient as they could be, suggesting that the attack would be even more powerful if they optimized their strategies. This motivates us to propose a few mitigations for address poisoning attacks at the protocol, contract, wallet, and user levels.
@@ 2.1 Blockchain primitives
Ethereum (and derived blockchains) features two types of addresses: wallet addresses (also called External Owned Accounts—EOA) and contract addresses. Wallet addresses to send Ether (ETH, the native currency) or interact with smart contracts (whose addresses are contract addresses). Smart contracts are executed by all nodes using the Ethereum Virtual Machine (EVM). Smart contracts are developed in a higher level programming language (e.g., Solidity), and then compiled to “bytecode” intelligible by the EVM before being deployed. Smart contracts allow users to perform actions, such as token implementation, token transfer. Of particular interest, ERC-20 tokens (e.g., USDT, USDC) are fungible assets, defined by the ERC-20 framework [40], for which creators specify a token name, symbol, and value unit. Essentially ERC-20 tokens allow developers to implement other cryptocurrencies or payment systems on top of Ethereum.
@@ 2.2 Wallet and blockchain address
(3) Hash (Keccak 256) the public key $K$, take the last 20 bytes, and add prefix $0x$ to obtain the address.
Elliptic curve (ECDSA) multiplication is computationally hard to invert, meaning that it is impractical to derive a private key from a public key. So, to generate a public key with a certain format (e.g., a vanity address), the only viable strategy is to perform a brute-force computation over potentially many private keys. The complexity of the brute force computation depends on the number of characters that have to be fixed in the desired public key address. For instance, it is roughly $16^{3}$ times harder to generate an (hexadecimal) address with a 3-character desired prefix compared to a random public key address. Attackers adopt this brute-force technique to generate lookalike addresses.
@@ 2.3 Attack characteristics
Blockchain address poisoning is markedly different from domain-name based phishing attacks in several respects. First, it focuses on the direct extraction of funds, and it may be more challenging for a victim to recover funds due to the irreversibility of cryptocurrency payments.
@@
Fourth, domain-name based attacks often proceed in multiple stages to steal funds: the victim first clicks on a malicious URL, is directed to a phishing website and is prompted to input sensitive information that can eventually be monetized. On the other hand, blockchain address poisoning is far more direct. Victims only need to make the single mistake of submitting transactions by selecting the attacker’s address.
Finally, blockchain address poisoning has differing costs to the attacker. They only need to deploy attack smart contracts (which can be even recycled from others for free), and one poisoning transfer costs only about a dollar in Ethereum and a cent in BSC. Generating lookalike addresses however has a variable cost that can be adjusted by the attacker, ranging anywhere from a laptop CPU to data centers of GPUs. Domain-based phishing attacks, on the other hand, require registering/hijacking domains, crafting phishing websites, and/or buying tool kits [3], all of which is expensive and time-intensive.
@@ 3.1 Threat model
We assume the existence of an attacker that passively monitors all public blockchains simultaneously in real-time. When available the attacker may also observe a state that has yet to be finalized in a chain such as transactions in the mempool. The attacker can also actively engage with any chain by launching contracts and sending transactions using many accounts, bounded only by the chain and economic constraints. Finally, the attacker can generate lookalike addresses at a cost consistent with (publicly known) state-of-the-art hardware.
We do not assume that the attacker has any out-of-band information about or agency towards a victim other than what can be obtained through public blockchains. Therefore attackers cannot influence the wallet software a victim is using, or make predictions about a victim’s transfers beyond what can be inferred through public signals. We also assume that the security models for the individual chains hold, so attackers cannot, for example, perform denial of service on a chain or drop/modify a victim’s transactions.
@@ 3.2 Transfer types
In a zero-value transfer (Fig. 1(b)), the attacker creates a transfer from $V$ to $L$ for the same token that $V$ sent to $R$, but with the value of zero. Typically, token transfers can only be sender-initiated. However, the current token standard features a public function called transferFrom, and popular token implementations allow the function callers to specify an arbitrary sender and an arbitrary receiver if the transferred value is zero. Therefore, the attacker can record a transfer event from $V$ to $L$ of the same ERC-20 token with zero value. Zero-value transfers are potentially even more dangerous than tiny transfers, as the victim appears to have previously sent money to the lookalike address.
@@ 3.3 Attack implementation
Figure 2 shows one method of implementing blockchain address poisoning in practice. Wallets are represented by rectangles and contracts (or ERC-20 tokens created by contracts) are represented by ovals. We color attackers’s instances in pink. First, an attack account ($A$) deploys a tailor-made attack contract ($AC$) and calls a function from it to execute multiple poisoning transfers at once. The attacker can also directly interact with the token contract to initiate a single poisoning transfer (without $AC$), which we observe in some cases despite its higher gas cost. The transaction $TX$ bundles together several potentially unrelated phishing transfers, in this case, the transfer of a lookalike token from $V1$ to $L1$ and a zero-value transfer from $V2$ to $L2$. Notice also that attacker $A1$ recycles their contract $AC1$ and $CT1$ to perform additional phishing transfers in $TX3$, and the infrastructure that $A1$ has embedded on a chain is typically distinct from that of other attackers such as $A2$’s use of $AC2$. It is also common for multiple attackers to target the same victim (e.g., $V1=V4$ ). As we will later describe, we have observed cases where attackers launch more than 100 poisoning transfers in a single transaction.
@@ 4.1 Detection algorithm
Ye et al. [46] identify zero-value transfers by enumerating all token transfers with zero value. While computationally efficient, this method does not consider tiny and counterfeit token transfers and possibly mislabel benign zero-value transfer attacks, e.g., those used for testing a destination address. Guan and Li [10] capture all three poisoning transfers but follow an approach slightly different from ours. They flag suspicious transfers based on attack characteristics first, whereas we filter transfers when a victim interacts with two improbably similar, but different addresses. Our approach appears more robust to attacker manipulation (refer to Appendix A). We deploy our own Ethereum and BSC full nodes to scan the data efficiently instead of relying on third-party RPC endpoints. Figure 3 sketches our detection system, which consists of five steps.
@@
We look at all transfers in stablecoins – i.e., we search all transfers sent to (or from) addresses in $\mathcal{V}$ for potential tiny (or zero value) transfers. We also retrieve all ERC-20 token transfers for which an address in $\mathcal{V}$ is a source for potential counterfeit token transfers. We denote the set of recently used addresses that interact with $\mathcal{V}$ by $\mathcal{C}=\{C_{1},C_{2},\ldots\}$.
@@
Step 4: Pruning possible attacks. For each victim $V$, the corresponding intended address $R$, and the associated lookalike address $L$, we extract all related token transfer events collected in step 2. In other words, we only look at the cases where the victim $V$ interacts with two very similar addresses $R$ and $L$, making false positives statistically unlikely (except for those caused by typing mistakes, which we separately discuss in Appendix C).
@@
      If a lookalike address $L$ sends an authentic token $AT$ to a victim $V$ with $0<v<v_{0}$ (tiny transfer attack). We set the threshold $v_{0}$ to 10 USD to exclude unreasonably high transfers from the attacker. [^6] More than 98.79% of tiny transfers below 10 USD are less than 3 USD. We manually found some rare cases where the attackers sent nearly 10 USD as a tiny transfer.
@@
  • A transfer is a payoff transfer if the victim $V$ sends an authentic token $AT$ ($v>0$) transfer to a lookalike address $L$. This is the transfer that ultimately monetizes the attack for the adversary.
The top-right table in Figure 3 summarizes our labeling scheme. Red cells are attackers. Note that we convert the transferred value into USD on the day of the transfer. The price data is from CoinGecko API based on a contract address.
@@
First, the victim could have legitimately interacted with two similar addresses by sheer coincidence. This is statistically unlikely, specifically because we already filter out addresses particularly vulnerable to the “birthday paradox” (refer to Appendix B for our detailed analysis).
@@
Third, the victim could have accidentally sent tokens to an incorrect destination address, e.g., due to a typing mistake. If 1) the destination address never sent any tokens, i.e., the attacker does not redeem the payoff transfer, and 2) $d$ is more than 20, i.e., computationally infeasible to generate such a $L$, we assume that the destination address was actually a typo. We indeed find 681 accidental transfers from 501 users, totaling 5.5 million USD in losses (almost all in Ethereum, with only 57,000 USD in BSC): Appendix C for more detailed method, results, and analysis.
@@ 4.2 Clustering attackers
We associate each poisoning transfer $TR$ with the corresponding attack instance $\{TX,L,CT,AC,A\}$. Similar to the “guilt-by-association” technique [43], we merge two attack sets if they share at least one of $TX$, $L$, or $A$. In other words, we assume that 1) two poisoning transfers in the same transaction belong to the same attacker, 2) an attacker is economically rational and specifies addresses $L$ they control (to successfully receive payoffs), and 3) an attacker does not share the private key of the attacker’s wallet $A$ with other attackers. We do not use information on $AC$ or $CT$ to avoid spurious cluster merges because anyone can call an attack contract or use a counterfeit token, regardless of source code availability. For instance, in Figure 2, we 1) merge ($TR_{1}$, $TR_{2}$) because both transfers are in the same transaction, 2) merge ($TR_{1}$, $TR_{3}$) if $L_{1}=L_{3}$, and 3) merge ($TR_{3}$, $TR_{4}$, $TR_{5}$, $TR_{6}$) if $A_{1}=A_{2}$.
Having applied this process to our entire dataset, we notice that two seemingly distinct large attack groups—active at different times or using different attack strategies—can be merged due to a few overlapping attack addresses $A$. Manual investigation revealed that those accounts appear to be bots copying poisoning transactions from two different attack groups with little or no modification, leading to a seemingly erroneous merge. Indeed, Guan and Li [10] performed similar clustering techniques, but did not consider the existence of bots and appeared to form one large cluster.
To prevent such issues, we explored several methods to exclude transactions from bots. While most of the approaches either still produce erroneous merges or remove too many transactions and underestimate the group size, analyzing account history effectively addresses both issues. For each attack address $A$, we additionally fetch their account history from Etherscan and calculate an attack ratio, defined as the proportion of transactions used for poisoning transfers. We then set a threshold to exclude addresses that have a low attack ratio. We document the details of those bots’ behavior and the bot exclusion process in Appendix D.

### v3

@@ Abstract
In many blockchains, e.g., Ethereum, Binance Smart Chain (BSC), the primary representation used for wallet addresses is a hardly memorable 40-digit hexadecimal string. As a result, users often select addresses from their recent transaction history, which enables blockchain address poisoning. The adversary first generates lookalike addresses similar to one with which the victim has previously interacted, and then engages with the victim to “poison” their transaction history. The goal is to have the victim mistakenly send tokens to the lookalike address, as opposed to the intended recipient. Compared to contemporary studies, this paper provides four notable contributions. First, we develop a detection system and perform measurements over two years on both Ethereum and BSC. We identify 13 times more attack attempts than reported previously—totaling 270M on-chain attacks targeting 17M victims. 6,633 incidents have caused at least 83.8M USD in losses, which makes blockchain address poisoning one of the largest cryptocurrency phishing schemes observed in the wild. Second, we analyze a few large attack entities using improved clustering techniques, and model attacker profitability and competition. Third, we reveal attack strategies—targeted populations, success conditions (address similarity, timing), and cross-chain attacks. Fourth, we mathematically define and simulate the lookalike address generation process across various software- and hardware-based implementations, and identify a large-scale attacker group that appears to use GPUs. We also discuss defensive countermeasures.
@@ 1 Introduction
Figure 1: Three types of poisoning transfers
@@
We design a detection system and run it between July 1, 2022, and June 30, 2024, on Ethereum and Binance Smart Chain (BSC). Combining both chains, we identify over 270M attack attempts (i.e., 13 times more than the previous efforts: 21M [47] and 14M [10]) that target over 17M victims. Out of those, attackers successfully received 6,633 transfers, adding up to over 83.8M USD in ill-gotten gains. This attack appears to be one of the largest cryptocurrency phishing schemes observed in the wild. We evaluate our detector by 1) the external dataset (100% precision, 97.2% recall), 2) analyzing its false negatives, 3) manually verifying the phished cases, 4) checking the smart contracts used for those attacks, and 5) comparing detection outcomes with different parameters (e.g., window size and similarity threshold).
Attackers typically optimize their operations by bundling several transfers into a single transaction and reusing addresses and contracts. This behavior allows us to link attack transfers through the “guilt-by-association” heuristic [44]. After removing transactions from unrelated bots that occasionally copy poisoning transactions from attackers, we can establish the existence of a few large distinct attack groups which we use to model the economics of address poisoning. For each group, we calculate revenues (i.e., a sum of phished amounts) and costs (i.e., total attack transaction costs). Despite high variability, most large groups manage to stay profitable by succeeding often, while smaller groups sometimes suffer from loss (by losing to the competition).
@@
Finally, we propose a few mitigations for address poisoning attacks at the protocol, contract, wallet, and user levels.
Compared to contemporary studies [47, 10], our paper offers the following contributions. First, our work detects attacks across multiple chains (BSC, Ethereum); in doing so, it identifies 13 times more attacks than prior work, and uncovers cross-chain attacks. Second, we devise improved clustering techniques, and use them to model group profitability and competition. Third, we perform novel experiments (e.g., victim profiling, success conditions, address generation modeling) to provide deeper insights into attack strategies and attacker computational capability.
@@ 2.1 Blockchain primitives
Ethereum (and derived blockchains) features two types of addresses: wallet addresses (also called External Owned Accounts—EOA) and contract addresses. Wallet addresses are to send Ether (ETH, the native currency) or interact with smart contracts (whose addresses are contract addresses). Smart contracts are executed by all nodes using the Ethereum Virtual Machine (EVM). Smart contracts are developed in a higher-level programming language (e.g., Solidity), and then compiled to “bytecode” intelligible by the EVM before being deployed. Smart contracts allow users to perform actions, such as token implementation, token transfer. Of particular interest, ERC-20 tokens (e.g., USDT, USDC) are fungible assets, defined by the ERC-20 framework [41], for which creators specify a token name, symbol, and value unit. Essentially, ERC-20 tokens allow developers to implement other cryptocurrencies or payment systems on top of Ethereum.
@@ 2.2 Wallet and blockchain address
(3) Hash (Keccak 256) the public key $K$, take the last 20 bytes, and add the prefix $0x$ to obtain the address.
Elliptic curve (ECDSA) multiplication is computationally hard to invert, meaning that it is impractical to derive a private key from a public key. To generate a public key with a certain format (e.g., a vanity address), the only viable strategy is to perform a brute-force computation over many private keys. The complexity of the brute force computation depends on the number of characters that have to be fixed in the desired public key address. For instance, it is roughly $16^{3}$ times harder to generate an (hexadecimal) address with a 3-character desired prefix compared to a random address. Attackers adopt this brute-force technique to generate lookalike addresses.
@@ 2.3 Attack characteristics
Blockchain address poisoning is markedly different from domain-name-based phishing attacks in several respects.
First, domain-name based attacks often proceed in multiple stages to steal funds: the victim first clicks on a malicious URL, is directed to a phishing website and is prompted to input sensitive information that can eventually be monetized [22]. On the other hand, blockchain address poisoning only requires a single mistake from the victim—erroneously selecting an attacker address instead of a legitimate one. Furthermore, fund recovery is far more challenging due to the irreversibility of cryptocurrency transactions.
@@
Finally, address poisoning has differing costs to the attacker. They only need to deploy attack smart contracts (which can be even recycled from others for free), and one poisoning transfer costs only about a dollar in Ethereum and a cent in BSC. Generating lookalike addresses however has a variable cost that can be adjusted by the attacker, ranging anywhere from a laptop CPU to data centers of GPUs. Domain-based phishing attacks, on the other hand, require registering/hijacking domains, crafting phishing websites, and/or buying tool kits [3].
@@ 3.1 Threat model
We assume the existence of an attacker that passively monitors all public blockchains simultaneously in real-time. When available the attacker may also observe a state that has yet to be finalized in a chain, such as transactions in the mempool. The attacker can also actively engage with any chain by launching contracts and sending transactions using many accounts, bounded only by the chain and economic constraints. Finally, the attacker can generate lookalike addresses at a cost consistent with (publicly known) state-of-the-art hardware.
We do not assume that the attacker has any out-of-band information about or agency towards a victim, other than what can be obtained through public blockchains. Therefore attackers cannot influence the wallet software a victim is using, or make predictions about a victim’s transfers beyond what can be inferred through public signals (i.e., they cannot tailor the attack to each specific victim). We also assume that the security models for the individual chains hold, so attackers cannot, for example, perform denial of service on a chain or drop/modify a victim’s transactions.
@@ 3.2 Transfer types
In a zero-value transfer (Fig. 1(b)), the attacker creates a transfer from $V$ to $L$ for the same token that $V$ sent to $R$, but with a value of zero. Typically, token transfers can only be sender-initiated. However, the current token standard features a public function called transferFrom, and popular token implementations allow the function callers to specify an arbitrary sender and an arbitrary receiver if the transferred value is zero. Therefore, the attacker can record a transfer event from $V$ to $L$ of the same ERC-20 token with zero value. Zero-value transfers are potentially even more dangerous than tiny transfers, as the victim appears to have previously sent money to the lookalike address.
@@ 3.3 Attack implementation
Figure 2 shows one method of implementing blockchain address poisoning in practice. Wallets are represented by rectangles and contracts (or ERC-20 tokens created by contracts) are represented by ovals. We color attackers’ instances in pink. First, an attack account ($A$) deploys a tailor-made attack contract ($AC$) and calls a function from it to execute multiple poisoning transfers at once. The attacker can also directly interact with the token contract to initiate a single poisoning transfer (without $AC$), which we observe in some cases despite its higher gas cost. The transaction $TX$ bundles together several potentially unrelated phishing transfers, in this case, the transfer of a counterfeit token from $V1$ to $L1$ and a zero-value transfer from $V2$ to $L2$. Notice also that attacker $A1$ recycles their contract $AC1$ and $CT1$ to perform $TX3$, and the infrastructure that $A1$ has embedded on a chain is typically distinct from that of other attackers, such as $A2$’s use of $AC2$. It is also common for multiple attackers to target the same victim (e.g., $V1=V4$ ), as we will later describe. We have observed cases where attackers launch more than 100 poisoning transfers in a single transaction.
@@ 4.1 Detection algorithm
Ye et al. [47] identify zero-value transfers by enumerating all token transfers with zero value. While computationally efficient, this method does not consider tiny and counterfeit token transfers and possibly mislabel benign zero-value transfer attacks, e.g., those used for testing a destination address. Guan and Li [10] capture all three poisoning transfers but follow an approach slightly different from ours. While both filter transfers when a victim interacts with two improbably similar, our approach appears more robust to attacker manipulation (refer to §5.2 and Appendix B). We deploy our own Ethereum and BSC full nodes to scan the data efficiently instead of relying on third-party RPC endpoints. Figure 3 sketches our detection system, which consists of five steps.
@@
We look at all transfers [^2] We capture all transfers that emit a “transfer” signature, including any contract actions. in stablecoins – i.e., we search all transfers sent to (or from) addresses in $\mathcal{V}$ for potential tiny (or zero value) transfers. We also retrieve all ERC-20 token transfers for which an address in $\mathcal{V}$ is a source for potential counterfeit token transfers. We denote the set of recently used addresses that interact with $\mathcal{V}$ by $\mathcal{C}=\{C_{1},C_{2},\ldots\}$.
@@
Step 4: Pruning possible attacks. For each victim $V$, the corresponding intended address $R$, and the associated lookalike address $L$, we extract all related token transfer events collected in step 2. In other words, we only look at the cases where the victim $V$ interacts with two very similar addresses $R$ and $L$, making false positives statistically unlikely (except for those caused by typing mistakes, which we separately discuss in Appendix D).
@@
      If a lookalike address $L$ sends an authentic token $AT$ to a victim $V$ with $0<v<v_{0}$ (tiny transfer attack). We set the threshold $v_{0}$ to 10 USD to exclude unreasonably high transfers from the attacker. [^7] While 98.79% of tiny transfers ($<10$ USD) are less than 3 USD, we manually confirmed some rare tiny transfers were almost 10 USD.
@@
  • A transfer is a payoff transfer if a victim $V$ sends an authentic token $AT$ ($v>0$) to a lookalike address $L$, which ultimately monetizes the attack for the adversary.
The top-right table in Figure 3 summarizes our labeling scheme. Pink cells are attackers. Note that we convert the transferred value into USD on the transfer date. The price data is from the CoinGecko API based on a contract address.
@@
First, the victim could have legitimately interacted with two similar addresses by sheer coincidence. This is statistically unlikely, specifically because we already filter out addresses particularly vulnerable to the “birthday paradox” (refer to Appendix C for our detailed analysis).
@@
Third, the victim could have accidentally sent tokens to an incorrect destination address, e.g., due to a typing mistake. If 1) the destination address never sent any tokens, i.e., the attacker does not redeem the payoff transfer, and 2) $d$ is more than 20, i.e., computationally infeasible to generate such a $L$, we assume that the destination address was actually a typo. We indeed find 681 accidental transfers from 501 users, totaling 5.5 million USD in losses (almost all in Ethereum, with only 57,000 USD in BSC): Appendix D for more detailed method, results, and analysis.
@@ 4.2 Clustering attackers
We associate each poisoning transfer $TR$ with the corresponding attack “transfer set” $\{TX,L,CT,AC,A\}$. Similar to the “guilt-by-association” technique [44], we merge two attack transfer sets if they share at least one of $TX$, $L$, or $A$. In other words, we assume that 1) two poisoning transfers in the same transaction belong to the same attacker, 2) an attacker is economically rational and specifies addresses $L$ they control (to successfully receive payoffs), and 3) an attacker does not share the private key of the attacker’s wallet $A$ with other attackers. We do not use information on $AC$ or $CT$ to avoid spurious cluster merges because anyone can call an attack contract or use a counterfeit token, regardless of source code availability. For instance, in Figure 2, we 1) merge ($TR_{1}$, $TR_{2}$) because both transfers are in the same transaction, 2) merge ($TR_{1}$, $TR_{3}$) if $L_{1}=L_{3}$, and 3) merge ($TR_{3}$, $TR_{4}$, $TR_{5}$, $TR_{6}$) if $A_{1}=A_{2}$. We keep merging attack transfer sets until there is no overlap of $TX$, $L$, and $A$ between attack groups.
Having applied this process to our entire dataset, we notice that two seemingly distinct large attack groups—active at different times or using different attack strategies—can be merged due to a few overlapping attack addresses $A$. Manual investigation revealed that those accounts appear to be bots copying poisoning transactions from two different attack groups with little or no modification, leading to a seemingly erroneous merge. Guan and Li [10] perform similar clustering techniques, but did not consider the existence of bots and appeared to form one large cluster.
To prevent such issues, we explored several methods to exclude transactions from bots. While most of the approaches either still produce erroneous merges or remove too many transactions and underestimate the group size, analyzing account history effectively addresses both issues. For each attack address $A$, we additionally fetch their account history from Etherscan and calculate an attack ratio, defined as the proportion of transactions used for poisoning transfers. We then set a threshold to exclude addresses that have a low attack ratio. We document the details of those bots’ behavior and the bot exclusion process in Appendix E.